### v2

# Ontological differentiation as a measure of semantic accuracy

Abstract
Understanding semantic relationships within complex networks derived from lexical resources is fundamental for network science and language modeling. While network embedding methods capture contextual similarity, quantifying semantic distance based directly on explicit definitional structure remains challenging. Accurate measures of semantic similarity allow for navigation on lexical networks based on maximizing semantic similarity in each navigation jump (Semantic Navigation, SN). This work introduces Ontological Differentiation (OD), a formal method for measuring divergence between concepts by analyzing overlap during recursive definition expansion. The methodology is applied to networks extracted from the Simple English Wiktionary, comparing OD scores with other measures of semantic similarity proposed in the literature (cosine similarity based on random-walk network exploration). We find weak correlations between direct pairwise OD scores and cosine similarities across $\sim$ 2 million word pairs, sampled from a pool representing over 50% of the entries in the Wiktionary lexicon. This establishes OD as a largely independent, definition-based semantic metric, whose orthogonality to cosine similarity becomes more pronounced when low-semantic-content terms were removed from the dataset. Additionally, we use cumulative OD scores to evaluate paths generated by vector-based SN and structurally optimal Shortest Paths (SP) across networks. We find SN paths consistently exhibit significantly lower cumulative OD scores than shortest paths, suggesting that SN produces trajectories more coherent with the dictionary’s definitional structure, as measured by OD. Ontological Differentiation thus provides a novel, definition-grounded tool for analyzing semantic structure and validating navigation processes in lexical networks.
I Introduction
The structure and dynamics of language represent a fascinating domain for complex systems science [1, 2, 3]. Lexical resources, such as dictionaries and thesauri, when represented as networks where words are nodes and definitional or synonymy relationships form links, exhibit intricate topologies reflecting the underlying semantic organization [4, 5, 6]. Understanding how meaning is encoded in these structures and how information propagates through them is crucial not only for linguistics [7], but also for advancing artificial intelligence, particularly in the development and validation of large language models and knowledge representations [8, 9].
A common approach to capture semantics involves embedding network nodes in vector spaces, often based on network connectivity patterns [10] or distributional statistics [11, 12]. For example, the Random Inheritance Method (RIM) generates such vectors by simulating random walks or information diffusion over the network [13, 14, 15], so that the vectors encode the probability of visiting each node (starting from every other node) after a fixed number of jumps of the associated Markov chain. Similarity between these feature vectors (e.g., cosine similarity) is then used as a proxy for semantic relatedness. This enables specific navigation strategies, a classic problem in the study of complex networks [16], such as the local-based semantic navigation approach proposed by Capitán et al. [17], which utilizes vector similarity to jump to the neighbor most similar in each navigation step, thus using Markovian dynamics on the network to guide paths between concepts. However, these embedding-based similarities and the paths they generate reflect statistical correlations or higher-order network proximity, rather than solely the explicit, declarative content of definitions [18]. Furthermore, standard graph-theoretical measures such as the shortest path distance often prioritize structural economy over semantic coherence, potentially traversing definitionally unrelated concepts. There remains a need for quantitative methods that directly utilize the intrinsic structure of definitions to measure semantic distance or path coherence, providing a complementary perspective to embedding-based approaches and a tool for evaluating semantic processes on these networks.
In this paper, we introduce and formalize Ontological Differentiation (OD), a method designed to quantify, based on first principles, the divergence between concepts based directly on their recursive definition structure within a lexical network. The core idea is that the semantic distance between two concepts relates to how quickly and extensively their definitional expansions overlap. We operationalize this by iteratively “reading” the definitions of a starting set of concepts and applying specific cancelation rules when elements are repeated across different expansion branches. The process continues until a termination condition is met, and the final OD score is calculated as a weighted sum reflecting the depth at which cancelations occurred. We focus primarily on Strong Ontological Differentiation (SOD), where cancelations are triggered only by overlaps between elements originating from different starting concepts.
Our primary goals are fourfold: (1) to provide a rigorous theoretical framework for defining OD as a meaningful measure of semantic proximity between terms; (2) to investigate the relationship between the definition-based SOD metric and the network-embedding-based cosine similarity derived from RIM; (3) to assess the robustness of these findings by applying our methods to networks constructed from the Simple English Wiktionary under three different definition processing schemes (ground-filtered, random removal, and targeted removal), simulating variations in corpus quality and authorship; and (4) having established the nature of the relationship between SOD and cosine similarity across these conditions, to apply cumulative SOD scores as an independent benchmark to evaluate the semantic coherence of paths generated by different navigation algorithms. Specifically, we use this framework to compare paths generated by Semantic Navigation (SN) [17], which optimizes RIM vector-based similarities, and standard Shortest Path (SP) navigation, which minimizes the overall path weight, using a large-scale sample of word pairs. By first understanding how SOD relates to cosine similarity, we can better interpret the results of using SOD to compare SN and SP paths.
Our results demonstrate that, while the precise correlation between SOD and RIM cosine similarity varies with the definition processing scheme (ranging from weak negative in less processed networks to negligible in the targeted removal network), SOD consistently acts as a largely independent measure of definitional content. Furthermore, and critically for its utility as an evaluation metric, we find that SN paths consistently exhibit lower cumulative SOD scores than SP paths across all tested network conditions. This suggests that SOD quantitatively validates the intuition that SN generates more definitionally coherent trajectories through the semantic network, a finding robust to significant variations in the underlying definitional structure.
The remainder of this paper is structured as follows: Section II details the theoretical framework of Ontological Differentiation and its variants. Section III describes the Semantic Navigation and Shortest Path algorithms used for comparison. Section IV outlines the three definition processing schemes leading to distinct datasets, network construction, vector computation, the SOD evaluation framework, the design of our large-scale experiment (including details on lexicon size and sampling), and an illustrative example. Section V first presents the comparative analysis of the relationship between direct SOD scores and RIM vector cosine similarity across the three conditions, and subsequently shows the results comparing the cumulative SOD scores of SN and SP paths. Finally, Section VI discusses the implications of our findings. Conceptual examples illustrating OD computation, together with detailed distributional plots for each dataset condition are provided in the Supplemental Material. In addition, the Supplemental Material includes a crucial extended analysis that investigates the confounding effects of network topology and path length, confirming the robustness of our central conclusions.
II First-principle definition of semantic similarity
The fundamental premise of our approach is that concepts within a structured lexicon, such as a dictionary, are constituted by their definitions—compositions of other concepts from the same lexicon. This inherently relational and compositional nature implies that the semantic “distance” or divergence between any two concepts, $P_{A}$ and $P_{B}$, can be understood as a measure of the definitional transformation required for their underlying structures to align. Specifically, we quantify this divergence by examining the extent and depth of recursive definitional expansion needed before the constituent elements of $P_{A}$ significantly overlap with those of $P_{B}$.
We formalize this through a framework termed Ontological Differentiation (OD). The “ontological” aspect reflects the view that this method addresses the relatedness of entities (here, lexical concepts) that are interdefined and exist through their compositional relationships. OD posits that, if concept $P_{A}$ is semantically proximal to $P_{B}$, their definitions, when iteratively expanded, will reveal shared constituent concepts at relatively shallow levels of recursion. Conversely, greater semantic distance implies that such convergence upon common definitional components will only manifest, if at all, after more extensive recursive unfolding, indicating a more substantial structural differentiation between them. OD operationalizes this by systematically tracking the repetition of elements—signifying shared definitional components—during this recursive expansion. The resulting divergence score is then a function of the depth at which these critical overlaps, which we name as cancelations, occur.
While the primary application and empirical validation in this paper focus on lexical semantics, the core OD methodology—quantifying divergence through the recursive depth needed for compositional alignment via specific cancelation and termination rules—is conceived as a generalizable framework. It holds potential for analyzing relationships in any system where entities are characterized by their interconnections and constituent parts, with the particular cancelation and termination rules adaptable to the specific nature of the entities and their relationships in the domain of interest.
The formal development of OD begins by representing a lexicon as a set of interdefined concepts:
Definition II.1.
Let $\mathbf{U}$ be a set of elements $\{P_{1},...,P_{n}\}$ of size $n\in\mathbb{N}$, for which every element $P_{i}\in\mathbf{U}$ is itself a set of size $m_{i}\in\mathbb{N}$ (can vary per element) of elements in $\mathbf{U}$, that is $P_{i}=\{P_{i_{1}},...,P_{i_{m_{i}}}\}$ where every $P_{i_{j}}\in\mathbf{U}$. This represents a collection of concepts where each concept is defined by a set of other concepts in the collection.
II.1 Ontological differentiation
We first define the process of expanding definitions recursively.
Definition II.2.
We call $R$ the read function, an operation such that for any $P_{i}\in\mathbf{U}$, $R(P_{i})$ yields the set $\{P_{i_{1}},...,P_{i_{m_{i}}}\}$ of elements in $\mathbf{U}$ constituting its definition. $R^{n}(P_{i})$ denotes the multiset of elements obtained by applying the read function recursively $n$ times.
Intuitively, this function “reads” the definition of term $P_{i}$ in the dictionary and outputs the set of words appearing in the definition. Formally, if we have as input some element $P_{i}\in\mathbf{U}$ whose content is given as $P_{i}=\{P_{i_{1}},...,P_{i_{m}}\}$, the read function $R^{1}(P_{i})$ will produce the multiset containing $P_{i_{1}},...,P_{i_{m}}$. The parameter $n$ specifies how many times this expansion operation is carried out, being $n=0$ just the reading $R^{0}(P_{i})=\{P_{i}\}$. For instance, $R^{2}(P_{i})$ represents the multiset formed by applying $R$ to each element in $R^{1}(P_{i})$. This can be generalized as $R^{n}(P_{i})=\bigcup_{P_{k}\in R^{n-1}(P_{i})}R(P_{k})$.
The read function can be applied not only to a single term in the dictionary, but to a subset of words. Then, the read function of the subset is naturally defined as the union of terms appearing in any definition of each element of the subset (i.e, the union of the output of the read function applied to each individual term in the subset).
Definition II.3.
Let $\mathcal{A}=\{P_{1},...,P_{k}\}\subseteq\mathbf{U}$ be an initial set of $k$ elements. Then, the read function $R^{\omega}(\mathcal{A})$ is defined as the collection of all elements generated by expanding each $P_{s}\in\mathcal{A}$ up to $\omega$ levels using the read function $R$. That is, $R^{\omega}(\mathcal{A})=\bigcup_{n=0}^{\omega}\bigcup_{s=1}^{k}R^{n}(P_{s})$.
The core idea of our definition of semantic proximity is related to the repetition of terms during this expansion process applied to the initial set $\mathcal{A}$. This is governed by specific rules. As discussed in the next subsection, different cancelation and termination rules produce different measures of semantic proximity.
Definition II.4.
A cancelation rule $\kappa$ determines the conditions under which an instance of an element $\alpha$ appearing in the expansion $R^{n}(P_{s})$ (originating from $P_{s}\in\mathcal{A}$ at level $n$) is marked as “canceled” or “repeated” based on its appearances elsewhere within the overall expansion $R^{\omega}(\mathcal{A})$.
Cancelation occurs when words arise repeatedly during the recursive expansion $R^{\omega}(\mathcal{A})$. As this expansion defines a hierarchy of levels (depending on the recursion level $n$), words will be more distant if repetitions arise higher in this hierarchy.
Definition II.5.
A termination rule $\tau$ determines the maximum level of expansion $\omega$, based on the state of canceled/uncanceled elements resulting from applying the cancelation rule $\kappa$.
The process of applying a specific cancelation rule $\kappa$ and termination rule $\tau$ to an initial set $\mathcal{A}$ is denoted $F_{\kappa\tau}(\mathcal{A})$. This process determines which elements get canceled at each level $n$ (from 00 to $\omega$) and the final expansion depth $\omega$. As a result, $F_{\kappa\tau}(\mathcal{A})$ yields a finite, recursive expansion of the terms included initially included in $\mathcal{A}$. We name the process $F_{\kappa\tau}(\mathcal{A})$ as the Ontological Differentiation of set $\mathcal{A}$.
Let $C_{nm}$ be the count of element instances originating from $P_{m}\in\mathcal{A}$ that are marked as canceled specifically at level $n$ according to rule $\kappa$. The above definitions allow us to quantify the divergence of this recursive definition expansion.
Definition II.6.
The score of the Ontological Differentiation process $F_{\kappa\tau}$ applied to $\mathcal{A}$ is given by the weighted sum of the number of canceled elements at each level:
$\text{Score}(F_{\kappa\tau}(\mathcal{A}))=\sum_{n=0}^{\omega}n\sum_{m=1}^{k}C_{nm}.$
This score quantifies the total “differentiation effort.” By weighting cancelations by the recursion level $n$ at which they occur, the score directly reflects our foundational premise: concepts are considered more distant (i.e., require a greater extent of definitional unfolding to find common ground) if their shared constituent elements—leading to cancelations—only emerge at deeper levels of their recursive expansion. Thus, the scoring method penalizes overlaps that occur at greater depths (higher $n$) and, conversely, gives less weight to (or “benefits”) those that occur at shallower levels (lower $n$), signifying closer semantic proximity. While this specific weighting scheme (multiplying by $n$) aligns with our focus on definitional depth in lexical networks, the general OD framework allows for this scoring function itself to be modified (e.g., using different weighting functions or incorporating other factors beyond depth) to suit the specific characteristics and desired interpretations of divergence in other domains or scenarios.
We finally define the overall distance between the elements in the set $\mathcal{A}$, which leads to a measure of semantic proximity between the terms in this set.
Definition II.7.
The Ontological Differentiation distance $\text{OD}(\mathcal{A})$ between the elements in the set $\mathcal{A}$ is defined as the score obtained from a specific differentiation process $F_{\kappa\tau}$: $\text{OD}(\mathcal{A})=\text{Score}(F_{\kappa\tau}(\mathcal{A}))$.
Naturally, the specific OD distance value depends on the chosen rules $\kappa$ and $\tau$, as well as the chosen scoring function.
It is important to note that the OD score, particularly the SOD variant used in our analysis, can grow very rapidly. The score is calculated as a sum of products, where one factor is the recursion level, $n$. As definitions are expanded recursively, both the number of potential cancellations and the recursion level $n$ increase, leading to a combinatorial growth in the potential score. Consequently, SOD scores for moderately distant concepts can easily span many orders of magnitude. This characteristic necessitates the use of logarithmic transformations for effective visualization and statistical analysis, as shown in our results.
II.2 Types of ontological differentiation
The general OD framework allows for various types, each defined by specific cancelation and termination rules. Here we introduce Strong Ontological Differentiation (SOD) first, due to its intuitive alignment with how semantic divergence might be perceived based on definitional content, including an illustrative example of its calculation based on dictionary terms. For that reason, in this work we focus primarily on strong ontological differentiation (SOD). However, in this subsection we briefly describe a second type of OD, named Weak Ontological Differentiation (WOD), to illustrate the flexibility and generality of our OD framework, as well as Great Ontological Differentiation (GOD), which is key in bounding the number of recursion steps in the expansion process. Conceptual examples further illustrating the computation of WOD and GOD scores, based on abstract concepts, are provided in the Supplemental Material.
II.2.1 Strong ontological differentiation (SOD)
Strong Ontological Differentiation (SOD) quantifies divergence by identifying elements that appear in expansions originating from different initial concepts. Repetitions occurring solely within the expansion of a single initial concept do not cause cancelation unless the element also appears in the expansion of another initial concept from the starting set $\mathcal{A}$. We formalize SOD cancelation and termination rules as follows:
Definition II.8.
(SOD cancelation rule, $\kappa_{S}$) An instance of element $\alpha$ generated at level $n$ from $P_{s}\in\mathcal{A}$ is canceled if $\alpha$ also appears in the expansion generated from a different initial element $P_{r}\in\mathcal{A}$ (where $r\neq s$) at any level $h\leq n$. Formally, if $\alpha\in R^{n}(P_{s})$ and there exists $\beta\in R^{h}(P_{r})$ such that $\alpha=\beta$ where $r\neq s$ and $h\leq n$, then $\alpha$ is canceled.
Definition II.9.
(SOD termination rule, $\tau_{S}$) The expansion stops at level $\omega$ if, after applying the cancelation rule $\kappa_{S}$ up to level $\omega$, all elements generated from at least one of the initial sets $P_{s}\in\mathcal{A}$ at level $\omega$ (i.e., all elements in the multiset $R^{\omega}(P_{s})$) have been canceled.
The specific design of these SOD rules is motivated by our aim to measure the definitional “distance” between two concepts $P_{s}$ and $P_{r}$ by tracking how one concept’s expanded definition progressively incorporates elements from the other. The $\kappa_{S}$ cancelation rule focuses exclusively on cross-side repetitions: an element is only canceled if it originates from one initial concept (e.g., $P_{s}$) and is also found within the expansion of the other initial concept ($P_{r}$). This distinction is crucial because we are interested in the convergence of initially distinct definitional paths; internal repetitions within a single concept’s expansion do not, by themselves, signify progress towards alignment with another concept. The $\tau_{S}$ termination rule is triggered when all elements comprising a specific expansion level of one initial concept (e.g., all words in $R^{1}(P_{s})$) are found to be “covered” or canceled by elements from the expansion of the other initial concept(s). This signifies a point where, at that particular depth of expansion, one concept’s definitional layer has been fully accounted for within the broader definitional structure of the other, indicating a significant degree of definitional subsumption or overlap.
We illustrate how SOD operates with the aforementioned rules by calculating the score for the pair of words ‘job’ and ‘business’. We will demonstrate the SOD calculation assuming that the definitions for these terms and their constituents are as provided in the “Definitions” column of Table II.2.1, as might be extracted from a lexical resource after a filtering process similar to that described later in Section IV.1.
The SOD process, detailed in Table II.2.1, unfolds by iteratively expanding these definitions. After each level of expansion, the SOD cancelation rule ($\kappa_{S}$) is applied by looking for elements appearing on both the ‘job’ and ‘business’ sides across all currently expanded levels. Subsequently, the SOD termination rule ($\tau_{S}$) is checked. In Table II.2.1, elements that are ultimately canceled by this process are shown in bold.
Definitions Level $\bm{P_{s}}=\mathrm{job}$ $\bm{P_{r}}=\mathrm{business}$ Scores
job: [pay, money, service] business: [buy, sell, service] pay: [give, money] money: [coin, note] service: [work, consume] buy: [pay, money, obtain] sell: [give, gain, money] Level 0 $R^{0}(\mathrm{job,business})$ job business $0\times 0=0$
Level 1 $R^{1}(\mathrm{job,business})$ pay money service buy sell service $4\times 1=4$
Level 2 $R^{2}(\mathrm{job,business})$ give money coin note work consume pay money obtain give gain money work consume $10\times 2=20$
Total distance = 24
Table 1: Example showing the calculation of the score SOD(job, business). The definitions of $P_{s}=$ ‘job’ and $P_{r}=$ ‘business’ include the terms ‘pay’, ‘money’, ‘service’, ‘buy’ and ‘sell’, which are expanded recursively according to their definitions (first column). The different multisets $R^{n}(P_{i})$, $i\in\{s,r\}$, obtained by recursive expansion at the $n$-th level of recursion are shown in columns $P_{s}$ and $P_{r}$, together with each level’s score (shown in the last column). Canceled elements are marked in boldface. In this example, after the second level expansion we observe that all the terms in the ‘job’ column at level 1 have appeared at levels less or equal than 2, so the process terminates at the second level. The score is calculated as the overall number of repetitions weighted by the level of recursion at which those repetitions occurred.
The step-by-step application of this SOD process to the (‘job’, ‘business’) pair unfolds as follows:
  • Level 0 ($R^{0}$):
    • $\circ$
      Expansion: ‘job’ and ’business’ are the initial elements.
    • $\circ$
      Cancelations: None, as no prior elements or cross-side repetitions exist.
    • $\circ$
      Termination check: No side is fully canceled at any level. The process continues.
  • Level 1 ($R^{1}$):
    • $\circ$
      Expansion: ‘job’ yields {pay, money, service}; ’business’ yields {buy, sell, service}.
    • $\circ$
      Cancelations: Considering levels L0-L1, we observe that the element ‘service’ appears in both $R^{1}(\text{job})$ and $R^{1}(\text{business})$. Thus, ‘service’ is marked for cancelation (in boldface) on both sides at level L1.
    • $\circ$
      Termination Check: No side at any level is fully canceled yet. The process continues to the next level of expansion.
  • Level 2 ($R^{2}$):
    • $\circ$
      Expansion: Elements from $R^{1}(\text{job})$ and $R^{1}(\text{business})$ are expanded according to the corresponding definitions (see Table II.2.1).
    • $\circ$
      Cancelations: Considering levels L0-L2, we recognize the following repetitions: (i) ‘give’, ‘money’, ‘work’, ‘consume’ appear in both $R^{2}(\text{job})$ and $R^{2}(\text{business})$; (ii) ‘pay’, from $R^{2}(\text{business})$, also appears in $R^{1}(\text{job})$; (iii) ‘money’, from $R^{2}(\text{business})$, also appears in $R^{1}(\text{job})$.
    • $\circ$
[…]
This example illustrates that cancelations in SOD are determined by cross-side repetitions considering all generated elements up to the current expansion depth. The process terminates when all elements originating from one of the initial concepts at a specific level of its expansion are canceled.
II.2.2 Weak ontological differentiation (WOD)
Weak Ontological Differentiation (WOD) offers a less restrictive approach to measuring divergence. Unlike SOD, which only considers cross-side repetitions for cancelation, WOD cancels an element if it has appeared anywhere before in the overall expansion $R^{\omega}(\mathcal{A})$, regardless of its origin side or level. WOD uses the following cancelation rule, while its termination rule ($\tau_{W}$) is the same as that defined for SOD ($\tau_{S}$).
Definition II.10.
(WOD cancelation rule, $\kappa_{W}$) An instance of element $\alpha$ generated at level $n$ from $P_{s}\in\mathcal{A}$ is canceled if $\alpha$ has appeared at least once before in the expansion $R^{\omega}(\mathcal{A})$ up to the current overall level $n_{c}\leq\omega$ being considered, regardless of its origin ($P_{r}$) or its own generation level ($h\leq n_{c}$). Formally, if $\alpha\in R^{n}(P_{s})$ (where $n\leq n_{c}$) and there exists $\beta\in R^{h}(P_{r})$ (where $h\leq n_{c}$) such that $\alpha=\beta$ for any $(h,r)\neq(n,s)$, then $\alpha$ is canceled.
This implies that, in WOD, any element appearing at least twice anywhere within the total expansion $R^{\omega}(\mathcal{A})$ will be marked as canceled (except for its first appearance). The process terminates, identically to SOD, when one side’s expansion at a particular level becomes entirely canceled. WOD is presented here primarily to illustrate that SOD is one specific instantiation within the broader OD framework; other OD types can be defined by varying the cancelation or termination rules.
A detailed, step-by-step illustration of WOD is provided in the Supplemental Material. This appendix example, like others provided for conceptual clarity, presents a more exhaustive, visual walkthrough of the iterative expansion and cancelation process. Such detailed tracking is excellent for understanding the core logic, but it is not the most computationally efficient representation for large-scale application. The actual calculations performed in this study utilize an optimized algorithm that tracks sets of uncanceled and canceled elements at each level for greater efficiency, as detailed and implemented in the code available in our repository [19].
II.2.3 Great ontological differentiation (GOD)
The Great Ontological Differentiation (GOD) process is not used for scoring divergence itself, but rather to determine a maximum relevant expansion depth, $\omega_{GOD}$. This is crucial for ensuring the finiteness of SOD and WOD calculations, especially in networks with cyclic definitions. GOD employs a distinct termination rule ($\tau_{G}$) based on the emergence of new unique elements during the expansion, which serve to determine an upper bound for the expansion depth.
Definition II.11.
(GOD termination rule, $\tau_{G}$) The expansion terminates at level $\omega$ if the set of all elements generated at level $\omega$, $\bigcup_{s=1}^{k}R^{\omega}(P_{s})$, contains no element $\alpha$ that was not already present in the union of elements from all preceding levels, $\bigcup_{n_{p}=0}^{\omega-1}\bigcup_{s=1}^{k}R^{n_{p}}(P_{s})$. We denote this termination level as $\omega_{GOD}$.
The level $\omega_{GOD}$, illustrated with an abstract example in the Supplemental Material, signifies the point beyond which no new concepts are introduced into the collective definitional expansion. Therefore, it serves as a practical upper bound for SOD/WOD: if the primary termination rule has not been met by $\omega_{GOD}+1$, the OD process can be halted. This prevents infinite loops and ensures the differentiation process has explored all potentially new definitional content before terminating due to stability rather than complete cancelation on one side.
II.2.4 Other types
The specific definitions chosen for the cancelation rule ($\kappa$) and the termination rule ($\tau$) dictate the behavior of any particular Ontological Differentiation process. The SOD, WOD, and GOD variants presented utilize straightforward rules based on element repetition, origin, and novelty. However, the general OD framework permits the formulation of numerous other rule sets. One could construct variants by modifying the existing conditions, combining elements of different rule types, introducing probabilistic elements, or defining entirely novel criteria for cancelation and termination based on different structural or sequential properties of the expansion. A full exploration of the mathematical possibilities afforded by altering these rules is beyond the scope of this work, where our aim is to introduce the core OD concept and demonstrate the utility of the SOD variant through empirical application.
III Navigation methods
In this section, we describe two distinct strategies for finding a route between a source node $s$ and a target node $t$ in a semantic network: (1) semantic navigation (SN), which leverages high‐order connectivity via vector embeddings, and (2) shortest‐path (SP) navigation, which finds cost‐minimal routes in the graph.
III.1 Graph representation
We model the semantic network as an undirected, weighted graph $G=(V,E)$, where: (i) $V=\{1,2,\dots,N\}$ is the set of $N$ nodes, each corresponding to a processed dictionary entry; and (ii) $E\subseteq V\times V$ is the set of edges.
The graph is constructed from the initial directed definition links. First, for each headword $i$ defined by a set of $m_{i}$ tokens, a directed edge is drawn from $i$ to each token $j$. The network is then symmetrized by adding a reverse edge $j\to i$ for every original edge $i\to j$. This choice is crucial for modeling the real, navigable space of a digital lexicon. For instance, if entry B is in the definition of A, a directed link $A\to B$ exists. However, A may not appear in the definition of B, meaning the reverse link $B\to A$ is absent. A user navigating from A to B can, however, always return to A (e.g., via a “back” button). To capture this fundamental bidirectional navigability, we enforce symmetry. This process reflects the principle that a definitional connection, even if unidirectional, implies a mutual semantic association that can be traversed in both directions. While this symmetric network forms the basis of our main results, we conduct a parallel analysis on the original (non-symmetrized) directed network in the Supplemental Material to validate our findings under stricter assumptions.
Each edge $(i,j)$ carries a non-negative, symmetric weight $w_{ij}=w_{ji}$. If the original directed link $i\to j$ existed because $j$ was one of the $m_{i}$ symbols in the definition of $i$, the weight is set to $w_{ij}=1/m_{i}$. The corresponding reverse edge $j\to i$ is assigned the same weight.
III.2 Semantic navigation
Semantic Navigation provides a method for finding paths between concepts by leveraging the relationships encoded in high-dimensional vector embeddings derived from the network structure. Specifically, we adapt the local-based semantic navigation strategy proposed by Capitán et al. [17], which operates as a greedy, cosine-similarity-driven walk using vectors generated via exploring network structure using random walks.
Semantic embeddings.
The Random Inheritance Method (RIM) [13] is a method inspired by exploring the network through random walks. A discrete-time Markov chain can be associated to the random walkers, whose transition probabilities are encoded in a row-stochastic transition matrix $P\in\mathbb{R}^{N\times N}$. In our implementation, this matrix reflects uniform transition probabilities based on node degrees in the symmetrized graph. Denoting by $\text{deg}(i)$ the degree of node $i$ in the symmetrized graph $G$, then the transition matrix is: $P_{ij}=1/\text{deg}(i)$ if $(i,j)\in E$ and $i\neq j$, and $P_{ij}=0$ otherwise. In such definition, we ignore self-loops. By construction, it is ensured that $\sum_{j}P_{ij}=1$. Note that the weights $w_{ij}$ are not used to construct $P$.
The elements of matrix powers $P^{k}$ encode the probabilities of jumping between pairs of nodes up to $k$ steps. This motivates the definition, for each node $i\in V$, of its semantic vector $\mathbf{v}_{i}\in\mathbb{R}^{N}$ as a vector encoding the random walk exploration of the network up to a certain inheritance depth level:
$\mathbf{v}_{i}\;=\;\sum_{k=1}^{T}\mathbf{e}_{i}P^{k},$
where (i) $\mathbf{e}_{i}$ is the $i$th standard basis vector in $\mathbb{R}^{N}$; (ii) $P^{k}$ is the $k$-step transition probability matrix derived from the Markov chain defined by the stochastic matrix $P$; (iii) $T$ is the inheritance depth, usually chosen to exceed the network’s diameter (we use $T=10$).
Thus $\mathbf{v}_{i}$ captures the cumulative reachability profile of node $i$ up to $T$ hops in the random network exploration defined by $P$. Pairs of semantic vectors $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$ are then used to measure the cosine similarity between nodes $i$ and $j$.
Cosine-based network navigation.
Given a source $s\in V$ and target $t\in V$ nodes in the network, SN constructs a path $\mathcal{P}=(p_{0},p_{1},\dots,p_{L})$ by iterating:
  1. 1.
    Initialization: Set $p_{0}=s$, step count $k=0$. Define a maximum path length $L_{\max}$ (it can be chosen as function of the graph diameter).
  2. 2.
    Iteration: While $p_{k}\neq t$ and $k<L_{\max}$:
    1. (a)
      Set current node $c=p_{k}$.
    2. (b)
      Let $\mathcal{N}(c)$ be the set of unvisited neighbors: $\mathcal{N}(c)=\{j\mid(c,j)\in E\text{ and }j\notin\{p_{0},\dots,p_{k}\}\}$.
    3. (c)
      If $\mathcal{N}(c)$ is empty, terminate (failure).
    4. (d)
      For each $j\in\mathcal{N}(c)$, compute the cosine similarity between $\mathbf{v}_{j}$ and $\mathbf{v}_{t}$:
      $\sigma_{j}\;=\;\frac{\mathbf{v}_{j}\cdot\mathbf{v}_{t}}{\|\mathbf{v}_{j}\|\|\mathbf{v}_{t}\|}$
      (handle potential zero vectors).
    5. (e)
      Choose the next node in the path as the one that maximizes cosine similarity with the current node $c$: $p_{k+1}=\arg\max_{j\in\mathcal{N}(c)}\sigma_{j}$.
    6. (f)
      Increment $k\leftarrow k+1$.
  3. 3.
    Termination: The algorithm stops when $p_{k}=t$ (success) or $k=L_{\max}$ or $\mathcal{N}(c)$ was empty. It returns a navigation path $\mathcal{P}=(p_{0},\dots,p_{k})$.
Because each move selects the neighbor whose embedding is most aligned with the target, SN aims to produce semantically coherent routes, albeit without guaranteeing minimal graph-theoretic cost.
III.3 Shortest‐path navigation
As a standard graph-theoretic baseline for comparison with SN, we consider shortest-path navigation. The objective of SP navigation is to find a path $\mathcal{Q}^{*}$ between a given source node $s$ and target node $t$ that minimizes the total cumulative edge weight. For any path $\mathcal{Q}=(q_{0}=s,q_{1},\dots,q_{L}=t)$ consisting of $L$ steps, this cumulative cost is defined as:
$\mathrm{Cost}(\mathcal{Q})\;=\;\sum_{\ell=1}^{L}w_{q_{\ell-1},q_{\ell}}$
where $w_{ij}$ are the non-negative, symmetric edge weights defined in Sec. IIIA.
It is important to note that the term “shortest path” in our study refers to this minimum-cost path, which is not necessarily the path with the fewest number of edges (hops). To compute this minimum-cost path $\mathcal{Q}^{*}$, we employ Dijkstra’s algorithm [20]. This widely used algorithm is guaranteed to find the path with the lowest cumulative cost from a single source node to all other nodes in a graph with non-negative edge weights. It operates by iteratively exploring the network outward from the source $s$, maintaining the lowest cost found so far to reach each node.
IV Evaluating Navigation Methods with SOD
To assess the utility of OD as a measure of semantic accuracy and to compare the pathways generated by Semantic Navigation and Shortest Path algorithms, we implemented a computational framework using data derived from the Simple English Wiktionary. This section details the data preparation, network construction, vector computation, evaluation methodology, and the design of the large-scale comparison experiment. Furthermore, an illustrative example is provided at the end of this section (Section IV.7) to demonstrate how cumulative SOD scores are used to distinguish between different navigation paths.
IV.1 Data acquisition and definition processing
Our primary data source was the Simple English Wiktionary (queried March 2025), chosen for its relatively controlled vocabulary and simpler definition structures. After an initial extraction and indexing process (detailed further in Sec. IV.6), we established a working lexicon of 19,283 unique entries for which definitions were available and which formed the basis for our subsequent experiments.
To assess the robustness of our findings and simulate variations in corpus characteristics (e.g., differing authorship, curation levels, or data sparsity) without introducing the many uncontrolled variables inherent in using entirely different external corpora, we generated three distinct definition datasets from this common 19,283-entry lexicon. This approach of systematically modifying a single source allows for a more controlled and theoretically grounded comparison of how variations in definitional structure impact the metrics. The generation of these datasets involved a foundational “ground filtering” stage applied to the definitions of these entries, followed by specific modifications to create each version.
Ground filtering (common to all datasets).
Initially, the definitions corresponding to our working lexicon underwent a common set of rigorous filtering and cleaning processes aimed at isolating core semantic relationships and standardizing the lexical base. For entries exhibiting polysemy, we adopted the simplifying convention of using only the first definition provided in the Wiktionary entry. This choice, while expedient for large-scale processing, acknowledges that not all facets of a word’s meaning may be captured. Entries’ definitions were then filtered based on their likely part-of-speech (only nouns, verbs, and adjectives retained) to focus on core lexical concepts. Furthermore, heuristics were applied to filter terms potentially representing inflected forms and to identify lemmas for definition tokens, aiming to build the network primarily from base forms. The text was then tokenized into constituent words. These tokens were further filtered: very short or numerical tokens were removed. Finally, to prevent potential trivial results in the subsequent OD calculations, any instance of a headword appearing within its own processed definition was removed.
Dataset 1: ground-filtered.
This dataset directly utilizes the definitions resulting from the “ground filtering” stage described above, with no further lexical modification of the definition content. It represents the most comprehensive but least specifically curated version of the definitions for the established lexicon and serves as the input for the other two processing schemes.
Dataset 2: random removal.
This dataset was created by taking the “ground-filtered” definitions and applying a structural perturbation. The ground-filtered definitions for the entries in our lexicon exhibit an average of 4.79 tokens per definition, with 2.32% of these definitions containing only a single token. For every definition in this set that contained more than one token (thus affecting 97.68% of the definitions), one token was randomly selected and removed. This process simulates the effect of data sparsity or less comprehensive definitions that might arise from different authorship or corpus limitations.
Dataset 3: targeted removal.
Starting from the “ground-filtered” definitions, this dataset underwent an additional, specific normalization step. A predefined list of 20 common, high-frequency, and often low-semantic-content words (analogous to stop-word removal) was applied to eliminate these terms from all definitions. Notably, all 20 of these removed words ranked within the top 100 most frequent tokens appearing in the ground-filtered definitions, and 13 of them were within the top 20. This aggressive, targeted filtering, detailed in our online repository [19] (which includes the list of removed words), was designed to create a more semantically focused network by removing potentially noisy functional or overly general terms.
These three distinct datasets, each based on the same lexicon but with varying definition content resulting from the ground-filtered, random removal, and targeted removal processing schemes, were then used independently for subsequent network construction, vector computation, and all comparative analyses.
IV.2 Network construction
For each of the three definition datasets (ground-filtered, random removal, and targeted removal), we independently constructed an undirected, weighted graph $G=(V,E)$ to represent the semantic relationships inherent in that specific dataset. The procedure for defining nodes, creating and symmetrizing edges, and assigning edge weights $w_{ij}$ followed the methodology detailed in Sec. IIIA. This crucial step resulted in three distinct network structures, providing the foundation for all subsequent pathfinding analyses.
IV.3 Semantic vector computation
To enable the semantic navigation algorithm, we computed semantic vectors $\mathbf{v}_{i}$ for each node within each of the three network structures generated in Sec. IV.2. Vectors were generated using RIM, following the procedure detailed in Sec. III.2. The specific implementation for each condition used the topology of the corresponding network to define the transition matrix $P$. This yielded three distinct sets of semantic vectors, ensuring that all SN-based analyses were performed using representations derived directly from the appropriate definitional dataset.
IV.4 Direct comparison of SOD and cosine similarity
Before using SOD to evaluate navigation paths, it is essential to understand its relationship to established vector-based semantic similarity measures. Therefore, a primary step in our analysis is to investigate the correlation between the proposed definition-based metric and a standard metric derived from network embeddings. Specifically, for a given pair of concepts represented by a source node $s\in V$ and a target node $t\in V$, we compute and compare two metrics:
  1. 1.
    The direct pairwise Strong Ontological Differentiation score, $\text{SOD}(s,t)$, calculated as described in Sec. II.
  2. 2.
    The cosine similarity, $\cos(\mathbf{v}_{s},\mathbf{v}_{t})$, calculated between the corresponding RIM vectors $\mathbf{v}_{s}$ and $\mathbf{v}_{t}$ (generated as described in Sec. IV.3, based on  [13, 17]).
Comparing these direct measures allows us to investigate whether the definition-based SOD metric captures similar or different information compared to the vector-embedding-based cosine similarity. This comparison is crucial for interpreting the path evaluation results presented in Sec. V. If SOD and cosine similarity were found to be highly correlated, then lower cumulative SOD scores for SN paths might simply reflect that SOD measures what SN optimizes (cosine similarity). Conversely, a lack of correlation would suggest SOD offers an independent validation based on definitional structure. Establishing this independence strengthens the claim that SOD measures a distinct facet of semantic relatedness, and that the observed alignment between SN paths and lower cumulative SOD is non-trivial. The computation of both direct metrics for each pair in our large-scale experiment facilitates this investigation.
IV.5 Path computation and SOD evaluation framework
Having established the context by comparing direct SOD and cosine similarity, our principal evaluation focuses on using cumulative SOD to quantitatively assess the semantic coherence of paths generated by the SN and SP navigation methods (detailed in Sec. III). For a given source-target pair $(s,t)$, we compute both the SN path $\mathcal{P}_{\rm SN}$ (via the greedy-cosine walk) and the SP path $\mathcal{Q}_{\rm SP}$ (via Dijkstra’s algorithm using weights $w_{ij}$).
To assess the overall semantic divergence of a path from its origin, we calculate a cumulative SOD score. This score reflects the total differentiation effort required to reach each node on the path from the starting node $p_{0}=s$. Specifically, for a path $\mathcal{P}=(p_{0},p_{1},\dots,p_{L})$, the cumulative SOD score, $\text{SOD}_{C}(\mathcal{P})$, is calculated as the sum of the pairwise SOD scores between the source node $p_{0}$ and every subsequent node $p_{i}$ ($i=1,\dots,L$) along the path:
$\text{SOD}_{C}(\mathcal{P})=\sum_{i=1}^{L}\text{SOD}(p_{0},p_{i})_{\omega_{GOD}+1}$
Here, $\omega_{GOD}+1$ is the maximum expansion depth for each pairwise calculation, which is determined beforehand by computing $\text{GOD}(p_{0},p_{i})$. This limit, indicated by the $\omega_{GOD}+1$ subscript, is imposed on the SOD calculation to guarantee termination and prevent potential infinite loops arising from cyclic definitions within the network. If the calculation reaches level $\omega=\omega_{GOD}+1$ without the termination rule being met, the process is halted at this level. This condition prevents infinite computation by ending that pair operation.
It is worth noting that alternative definitions for a path’s cumulative SOD score exist. For instance, one could sum the SOD scores between each intermediate node $p_{i}$ ($i=0,\dots,L-1$) and the final target node $p_{L}=t$. While potentially valid, we selected the current formulation, $\sum_{i=1}^{L}\text{SOD}(p_{0},p_{i})$, as it directly measures the accumulated definitional divergence from the origin along the path, which aligns conceptually with assessing how well a path maintains coherence relative to its starting point.
Another alternative, summing the SOD scores between consecutive nodes, $\sum_{i=0}^{L-1}\text{SOD}(p_{i},p_{i+1})$, was considered less suitable for differentiating between SN and SP paths. Given that navigation steps often involve moving between closely related concepts (especially in SN, but also often in SP due to definition links), the pairwise $\text{SOD}(p_{i},p_{i+1})$ is frequently very low (e.g., 1 or 2 if $p_{i+1}$ is directly in the definition of $p_{i}$). Summing these minimal step-wise scores might result in path scores that are too similar or dominated by path length rather than semantic jumps, potentially obscuring differences between the navigation strategies.
Furthermore, our chosen definition explicitly includes the term $\text{SOD}(p_{0},p_{L})$, representing the direct definitional distance between the source and the target. While omitting this final term is conceivable, its inclusion ensures that the cumulative score reflects the overall divergence of the entire source-target pair, providing a baseline dimension against which the path’s intermediate divergence is accumulated.
Finally, regarding the interpretation of the metric used in the Results Section (Sec. V), a lower value of $\text{SOD}_{C}(\mathcal{P})$ indicates that the sequence of nodes forming the path $\mathcal{P}$ exhibits less overall definitional divergence from the origin $p_{0}$. Consequently, paths with lower scores are considered more definitionally coherent or semantically proximal to the source according to the SOD framework.
IV.6 Large-scale experiment design
To perform a robust comparison between SN and SP using the cumulative SOD metric, and to analyze the relationship between direct SOD and cosine similarity across the different definition processing schemes (as discussed in Sec. IV.4), we conducted a large-scale computational experiment replicated across each of our three dataset conditions (ground-filtered, random removal, targeted removal).
The core experimental design involved generating a large, consistent set of source-target pairs and then computing all relevant metrics for these pairs within the context of each specific dataset condition.
Source-target pair generation.
We defined a sampling strategy common to all conditions:
  1. 1.
    Fixed set: A curated set of 100 source concepts was selected, aiming to cover a diverse range of semantic domains (e.g., abstract concepts, living beings, objects; see [19] for the list). All 100 fixed words were confirmed to be present in our working lexicon.
  2. 2.
    Random set pool: The pool of potential random concepts consisted of the 19,283 unique entries identified after the initial indexing and ground filtering stages (corresponding to the headwords in the ground-filtered dataset, described in Sec. IV.1).
  3. 3.
    Random set selection: A large sample of 10,000 target concepts was randomly selected without replacement from this 19,283-entry pool. This sample represents $51.86\%$ of the eligible lexicon in this range, providing substantial coverage.
  4. 4.
    Pair generation: Pairs were formed by considering both directions between the 100 fixed words and the 10,000 selected random words, i.e., (fixed, random) and (random, fixed). This resulted in approximately 2 million distinct source-target pairs used for evaluation in each experimental condition.
Metric computation for each condition.
Crucially, the subsequent computation of paths and metrics was performed independently for each of the three dataset conditions. For each generated pair $(s,t)$, and for each condition (ground-filtered, random removal, targeted removal), we computed:
  1. 1.
    The SN path $\mathcal{P}_{\rm SN}$ and the SP path $\mathcal{Q}_{\rm SP}$, using the specific network graph $G$ and edge weights $w_{ij}$ derived from that condition’s definitions (Sec. IV.2). SN pathfinding utilized the RIM vectors $\mathbf{v}_{i}$ specific to that condition (Sec. IV.3).
  2. 2.
    The corresponding cumulative SOD scores, $\text{SOD}_{C}(\mathcal{P}_{\rm SN})$ and $\text{SOD}_{C}(\mathcal{Q}_{\rm SP})$, using the SOD framework applied to that condition’s definition set (Sec. IV.5).
  3. 3.
    The direct pairwise score $\text{SOD}(s,t)$, again using the definitions from the specific condition.
  4. 4.
    The direct cosine similarity $\cos(\mathbf{v}_{s},\mathbf{v}_{t})$, using the RIM vectors $\mathbf{v}_{s}$ and $\mathbf{v}_{t}$ specific to that condition.
This comprehensive process resulted in three parallel datasets, each containing the $\sim 2$ million pairs along with their associated navigation paths and SOD/cosine metrics calculated relative to the specific definition processing scheme (ground-filtered, random removal, or targeted removal). These datasets form the basis for the comparative statistical analyses presented in Sec. V.
IV.7 Illustrative example
To further clarify the SOD calculation and its application to path evaluation, we provide an example using the definitions and network structure from the targeted removal dataset condition (Sec. IV.1).
IV.7.1 Example: path comparison
To illustrate how cumulative SOD distinguishes between navigation strategies, consider the paths generated between the source node ‘zebra’ and the target node ‘circle’. Using SN, which follows maximal cosine similarity to the target vector at each step, the path generated was:
[…]
Using SP navigation based on Dijkstra’s algorithm with edge weights $w_{ij}=1/m_{i}$, the path found was:
[…]
We then calculated the cumulative SOD score for each path using the formula $\text{SOD}_{C}(\mathcal{P})=\sum_{i=1}^{L}\text{SOD}(\text{origin},p_{i})$. The pairwise SOD scores between the origin (‘zebra’) and each subsequent node on the paths were computed (determining $\omega_{GOD}+1$ for each pair): SN path contributions:
SOD(zebra, animal) = 1
SOD(zebra, move) = 7
SOD(zebra, whirl) = 40385
SOD(zebra, circle) = 4659
SP path contributions:
SOD(zebra, looks) = 1
SOD(zebra, full moon) = 388098
SOD(zebra, circle) = 4659
Summing these contributions yields the total cumulative SOD scores, $\text{SOD}_{C}(\mathcal{P}_{\rm SN})=1+7+40385+4659=45052$ and $\text{SOD}_{C}(\mathcal{Q}_{\rm SP})=1+388098+4659=392758$.
[…]
Conversely, the SN path avoids such large definitional jumps relative to the origin, featuring intermediate concepts (‘animal’, ‘move’, ‘whirl’) that maintain closer definitional relatedness to ‘zebra’. The cumulative SOD metric quantitatively captures this difference, assigning a lower overall divergence score to the SN path. This example thus illustrates our central hypothesis: the cumulative SOD score, derived purely from definitional structure and independent of the vector space, serves to formalize semantic intuition about path coherence. It provides a quantitative basis for observing that navigation guided by semantic embeddings (SN) tends to produce paths with lower definitional divergence compared to structurally shortest paths (SP), a pattern whose generality we investigate across the large-scale dataset in the following section.
V Results
Using the methodology outlined in Sec. IV, involving three distinct definition processing schemes (ground-filtered, random removal, targeted removal), we conducted a large-scale computational analysis on the Simple English Wiktionary network. For each condition, approximately 2 million source-target word pairs were evaluated. The results presented below compare the metrics across these three conditions. We first analyze the relationship between direct SOD scores and RIM cosine similarity, followed by the comparison of cumulative SOD scores for Semantic Navigation (SN) versus Shortest Path (SP) navigation paths. Note that due to path generation failures for some pairs or SOD calculations terminating early based on the $\omega_{GOD}+1$ termination condition for others, the exact number of valid pairs considered may vary slightly between analyses and conditions, but all are approximately $N\approx 2.0\times 10^{6}$.
V.1 Relationship between direct SOD and cosine similarity
Our first analysis investigates the relationship between our proposed definition-based distance metric, the direct pairwise score $\text{SOD}(s,t)$, and a standard measure of semantic similarity derived from network structure, the cosine similarity $\cos(\mathbf{v}_{s},\mathbf{v}_{t})$ between the RIM vectors. This was performed comparatively across the three dataset conditions.
Distributions and transformations.
The raw distributions of the two metrics differ substantially and vary across conditions. Detailed plots showing the distributions at each transformation stage (raw, log(1+x), normalized log(1+x)) for each condition are provided in the Supplementary Material (Figs. S1–S3). Consistently across all conditions, the raw SOD scores exhibit an extremely heavy right tail (see Figs. S1A, S2A, S3A), spanning many orders of magnitude and necessitating transformation for meaningful analysis. Raw cosine similarity values are bounded but also show skewness towards lower values (Figs. S1D, S2D, S3D). Given the extreme skewness, particularly of SOD, we applied a $\log(1+x)$ transformation to both metrics, a standard practice for handling such data. Figure 1 presents comparative box plots of the resulting log(1+SOD) and log(1+Cosine) distributions across the three conditions. These plots reveal differences in the central tendency and spread; for instance, the median log(1+Cosine) is notably lower for the targeted removal condition compared to the other two, while the log(1+SOD) distributions show variations primarily in their interquartile range and the extent of outliers. Kolmogorov-Smirnov tests confirmed that these log-transformed distributions differ significantly between all pairs of conditions (p $<$ 0.0001; see Sec. V.3), underscoring that the definition processing significantly impacts the resulting metric distributions.
Figure 1: Comparative box plots showing the distributions of (a) log(1+SOD) scores and (b) log(1+Cosine) similarity (bottom) across the three dataset conditions: ground-filtered, random removal, and targeted removal. The plots illustrate differences in median (orange line), interquartile range (box), and outlier extent (whiskers and points).
Visual correlation.
The relationship between the log-transformed metrics is visualized using comparative 2D density histograms in Fig. 2. Across all three conditions, the plots show dense clouds concentrated at lower log(1+SOD) values, with no discernible linear or non-linear correlation structure. The density appears most concentrated (less diffuse) in the targeted removal condition compared to the ground-filtered and random removal conditions. This visual inspection strongly suggests a weak or negligible correlation between the log-transformed SOD and cosine similarity metrics, regardless of the definition processing scheme.
Figure 2: Comparative 2D density histograms showing the joint distribution of log(1+Cosine) vs. log(1+SOD) for the ground-filtered (a), random removal (b), and targeted removal (c) conditions. Color intensity represents the count of word pairs within each bin on a logarithmic scale. The lack of a clear trend or diagonal structure in all panels visually confirms the weak correlation between the two log-transformed metrics across all conditions.
Rank correlation.
To quantify the monotonic relationship between the original metrics, we calculated the Spearman rank correlation coefficient ($\rho$) between the raw $\text{SOD}(s,t)$ scores and the raw $\cos(\mathbf{v}_{s},\mathbf{v}_{t})$ values for each condition. Table 2 summarizes these results. A clear trend emerges: the ground-filtered dataset exhibits a weak negative correlation ($\rho\approx-0.21$), which diminishes for the random removal dataset ($\rho\approx-0.10$) and becomes practically negligible for the targeted removal dataset ($\rho\approx-0.03$). While statistically significant due to the large sample sizes ($p<0.0001$ for all three conditions), even the strongest correlation observed ($-0.21$) is weak in practical terms. This indicates no meaningful rank-order relationship in the targeted removal case, and only a very slight tendency for higher raw SOD scores to correspond with lower raw cosine similarities in the less processed conditions.
Condition Valid Pairs Spearman’s $\rho$ (p-value)
Ground-filtered $\approx 2.0\times 10^{6}$0.2069 ($p<0.0001$)
Random removal $\approx 2.0\times 10^{6}$0.0984 ($p<0.0001$)
Targeted removal $\approx 2.0\times 10^{6}$0.0305 ($p<0.0001$)
Table 2: Comparative Spearman rank correlation (raw SOD vs. raw cosine similarity) across conditions.
Percentile agreement.
We further assessed agreement by comparing the relative rankings of pairs after $\log(1+x)$ transformation and min–max normalization, dividing the $[0,1]$ range into 10 decile bins. Figure 3 visually compares the normalized distributions of log(1+SOD) and log(1+Cosine) for each condition, highlighting the general mismatch in shape and peak location between the two metrics within each condition. Table 3 presents the percentage of pairs falling into the same decile bin for both metrics. Overall agreement is low across all conditions, confirming that the relative ordering of pairs by SOD differs substantially from the ordering by cosine similarity. Interestingly, the agreement is highest for the ground-filtered dataset (17.73%), followed closely by random removal (17.37%), and is lowest for the targeted removal dataset (15.07%).
Figure 3: Comparison of the probability density functions for normalized log(1+SOD) and normalized log(1+Cosine) for the ground-filtered (a), random removal (b), and targeted removal (c) conditions. The distinct shapes and peak locations within each panel visually confirm the generally low percentile agreement found between the two metrics’ relative rankings (quantified in Table 3).
Condition Valid Pairs Overall Agreement (%)
Ground-filtered $\approx 2.0\times 10^{6}$ 17.73%
Random removal $\approx 2.0\times 10^{6}$ 17.37%
Targeted removal $\approx 2.0\times 10^{6}$ 15.07%
Table 3: Comparative percentile agreement analysis: normalized log(1+SOD) vs. normalized log(1+Cosine) (10 Bins).
Summary of SOD vs. cosine comparison.
Taken together, these analyses consistently demonstrate a weak relationship between direct SOD scores and RIM cosine similarity across all dataset conditions. This relationship, characterized generally by weak negative rank correlation (on raw data) and low percentile agreement (after log transformation), becomes progressively weaker, approaching statistical orthogonality, as definitions undergo more specific curation (from ground-filtered to targeted removal). This trend strongly suggests that SOD captures distinct aspects of semantic relatedness grounded in explicit definitional structure, which are largely independent of the network proximity information captured by RIM cosine similarity, particularly when applied to a more curated, semantically focused network like the targeted removal dataset.
V.2 Comparison of cumulative SOD for SN vs. SP paths
Next, we evaluated the semantic coherence of paths generated by the SN and SP algorithms by comparing their cumulative SOD scores. For a given path $\mathcal{P}$, this score is defined as $\text{SOD}_{C}(\mathcal{P})=\sum_{i=1}^{L}\text{SOD}(p_{0},p_{i})_{\omega_{GOD}+1}$ (see Sec. IV.5). This evaluation was performed across the  2 million source-target pairs for each of the three dataset conditions. A lower $\text{SOD}_{C}$ score indicates greater definitional coherence relative to the path origin.
Condition SN $<$ SP SN $=$ SP SN $>$ SP
Ground-filtered 59.82% 16.96% 23.22%
Random removal 62.42% 14.80% 22.78%
Targeted removal 58.22% 15.48% 26.30%
Table 4: Comparison of path outcomes based on cumulative SOD scores ($\text{SOD}_{C}$) for semantic navigation (SN) vs. shortest path (SP) across conditions. Headers indicate the comparison of $\text{SOD}_{C}(\text{SN})$ relative to $\text{SOD}_{C}(\text{SP})$.
Dominance of semantic navigation.
Table 4 presents the aggregated results comparing the $\text{SOD}_{C}$ scores for paths generated by SN versus those by SP between the same endpoints for each condition. A consistent and robust finding emerges across all three processing schemes: paths generated by SN exhibit lower $\text{SOD}_{C}$ scores than paths generated by SP in a significant majority of cases (59.82% for ground-filtered, 62.42% for random removal, and 58.22% for targeted removal). This primary observation holds regardless of the definition processing method applied. Statistical tests further confirm the significance of this dominance within each condition: $\chi^{2}$ goodness-of-fit tests strongly reject the hypothesis of equal likelihood for the three outcomes, namely: $\text{SOD}_{C}(\text{SN})<\text{SOD}_{C}(\text{SP})$, $\text{SOD}_{C}(\text{SN})=\text{SOD}_{C}(\text{SP})$, $\text{SOD}_{C}(\text{SN})>\text{SOD}_{C}(\text{SP})$ ($p<$ 0.0001 for all conditions). Furthermore, one-sided binomial tests confirm that the proportion of pairs where $\text{SOD}_{C}(\text{SN})<\text{SOD}_{C}(\text{SP})$ is significantly greater than 50% for each condition ($p<$ 0.0001). This provides strong evidence that SN consistently generates paths more definitionally coherent with the origin (as measured by lower $\text{SOD}_{C}$) compared to SP, demonstrating the robustness of using this cumulative SOD metric for validating semantic navigation across different corpus characteristics. This fundamental result is further tested and confirmed on more restrictive directed networks in the Supplemental Material, where we also control for the confounding effect of path length.
Sensitivity of comparison outcomes to network conditions.
While the core finding—that SN paths are generally favored by the $\text{SOD}_{C}$ metric—is robust, the precise distribution of the three comparison outcomes, $\text{SOD}_{C}(\text{SN})$ relative to $\text{SOD}_{C}(\text{SP})$, does show statistically significant variation across the different network conditions. This is confirmed by a $\chi^{2}$ test for contingency applied to the counts underlying Table 4 ($\chi^{2}=12156.70$, $\text{df}=4$, $p<0.0001$), indicating that the definition processing scheme influences the quantitative results of the SN vs. SP comparison. Notably, the random removal condition shows the highest proportion of cases where $\text{SOD}_{C}(\text{SN})<\text{SOD}_{C}(\text{SP})$ (62.42%). Conversely, the targeted removal condition exhibits the lowest proportion in this category (58.22%) and the highest proportion where $\text{SOD}_{C}(\text{SN})>\text{SOD}_{C}(\text{SP})$ (26.30%). This suggests that the specific structural changes induced by different processing methods (e.g., random noise vs. targeted removal of common words) can subtly alter the relative performance of SN and SP algorithms when evaluated using the definition-based $\text{SOD}_{C}$ score. For instance, removing random links might disproportionately affect the structurally optimal SP paths, enhancing SN’s relative advantage, whereas targeted removal might occasionally create shortcuts usable by SP that are also definitionally sound. These variations highlight the interplay between network structure and navigation algorithm performance, but they do not undermine the consistent overall observation that SN paths align better with definitional coherence as measured by cumulative SOD.
V.3 Statistical comparison of metric distributions
To further quantify the impact of the definition processing schemes, we performed two-sample Kolmogorov-Smirnov (KS) tests comparing the distributions of the log-transformed metrics between pairs of conditions. The results (Table 5) show statistically significant differences ($p<$ 0.0001) for both log(1+SOD) and log(1+Cosine) distributions between all pairs of conditions (ground-filtered vs. random removal, ground-filtered vs. targeted removal, and random removal vs. targeted removal). These statistical differences align with the visual distinctions observed in the comparative box plots (Fig. 1). This confirms that the different definition processing strategies indeed result in measurably distinct statistical properties for both the SOD and cosine similarity metrics derived from the resulting networks.
Comparison Metric KS Stat (p-value)
GF vs RR log(1+SOD) 0.1842 ($p\ll 0.0001$)
log(1+Cosine) 0.3216 ($p\ll 0.0001$)
GF vs TR log(1+SOD) 0.1006 ($p\ll 0.0001$)
log(1+Cosine) 0.6760 ($p\ll 0.0001$)
RR vs TR log(1+SOD) 0.0864 ($p\ll 0.0001$)
log(1+Cosine) 0.4088 ($p\ll 0.0001$)
Table 5: Kolmogorov-Smirnov (KS) 2-sample test results comparing log(1+SOD) and log(1+Cosine) distributions between conditions: ground-filtered (GF), random removal (RR), targeted removal (TR).
VI Conclusions
Quantifying semantic relationships directly from the structure of lexical networks remains a fundamental challenge. While vector embedding techniques capture contextual proximity, intrinsic measures reflecting explicit definitional relationships are less explored. In this work, we introduced and formalized Ontological Differentiation (OD), specifically Strong Ontological Differentiation (SOD), as a novel method to compute a divergence score based purely on the recursive expansion and overlap of their definitions. We aimed to demonstrate its utility as a measure of semantic structure and coherence within these networks. To establish the robustness and general applicability of our findings beyond the specifics of any single corpus representation, we performed our analyses comparatively across three variations of a network derived from the Simple English Wiktionary, created using distinct definition processing schemes (ground-filtered, random removal, and targeted removal) to simulate diverse corpus characteristics.
Our investigation yielded two primary conclusions. First, we examined the relationship between direct pairwise SOD scores and RIM cosine similarity across the three network conditions. The results revealed that this relationship is sensitive to definition processing: a weak negative correlation in the ground-filtered network ($\rho\approx-0.21$) progressively diminished to become negligible in the targeted removal network ($\rho\approx-0.03$) (Table 2). This trend, supported by percentile agreement analyses (Table 3), establishes that SOD captures aspects of semantic relatedness grounded in explicit definitions that are largely orthogonal to embedding-based similarity, particularly as the definitional structure is refined. This confirms SOD offers a complementary perspective to standard similarity measures.
Second, we utilized the cumulative SOD score ($\text{SOD}_{C}$) as an independent benchmark to evaluate the semantic coherence of paths generated by semantic navigation versus shortest path algorithms. This core evaluation, replicated across all three network conditions, demonstrated the utility and robustness of the $\text{SOD}_{C}$ metric. We found that SN paths consistently exhibited significantly lower $\text{SOD}_{C}$ scores compared to SP paths across all three datasets (Table 4), indicating greater definitional coherence relative to the origin. This finding was proved to be statistically robust across all three network conditions —whether the definitions were minimally processed, randomly perturbed, or specifically curated— with one-sided binomial tests for each confirming that SN paths outperform SP paths significantly more often than the 50% chance level ($p<0.0001$). The consistency of this result across the simulated variations in corpus structure provides strong evidence that cumulative SOD serves as a generally effective metric for validating the semantic coherence of navigation processes relative to definitional structure. While the precise magnitude of SN’s advantage varied significantly between conditions (as shown by a $\chi^{2}$ test, $p<$ 0.0001), its overall superiority as measured by $\text{SOD}_{C}$ remained constant.
In essence, Ontological Differentiation provides a formal method grounded in shared definitional components. Our comparative analysis across systematically varied definition sets validates SOD as a definition-based measure of semantic divergence, distinct from vector similarity. Most importantly, its consistent ability to identify SN paths as more definitionally coherent than SP paths across varied network structures demonstrates its robustness and utility as a tool for analyzing and validating semantic processes in lexical networks, offering insights independent of specific corpus construction details. This conclusion is not only supported by our main analysis on symmetric networks but is decisively reinforced by our extended analysis in the Supplemental Material. There, we demonstrate that SN’s superiority holds even on structurally constrained directed networks once the confounding effect of path length is properly controlled, confirming that the semantic coherence captured by SN is a fundamental advantage of the navigation strategy itself, independent of specific corpus construction or network topology.
Acknowledgments
Funding was provided by the PRIORITY grant (PID2021-127202NB-C22) to JAC, and has been partially supported by the grant PID2021-122711NB-C21, both funded by MCIN / AEI / 10.13039 / 501100011033 and “ERDF A way of making Europe”.
The authors acknowledge computing resources at the Magerit Supercomputer of the Universidad Politécnica de Madrid.
Supplemental Material for
Ontological differentiation as a measure of semantic accuracy P. García-Cuadrillero, F. Revuelta and J. A. Capitán
This Supplemental Material provides additional details to support the main text. It is organized into three sections:
  1. 1.
    Conceptual examples: We first present a series of simplified, abstract examples to illustrate the core mechanics of the different Ontological Differentiation (OD) variants, including Weak Ontological Differentiation (WOD) and Great Ontological Differentiation (GOD).
  2. 2.
    Metric distributions: We then provide detailed visualizations of the distributions for Strong Ontological Differentiation (SOD) scores and RIM cosine similarity values. These plots show the data at different transformation stages for each of the three dataset conditions analyzed in the main paper: ground-filtered, random remocal, and targeted removal.
  3. 3.
    Extended analysis: To conduct a stricter test of semantic coherence, we extended the main text analysis using a non-symmetrized, directed network, in which the links strictly follow the definitional relationships arising in the datasets.
Appendix A Conceptual Examples of OD construction
A.1 Sample Definition Set
The following conceptual examples utilize a small, illustrative set of abstract concepts and their definitions (a “sample set”) to demonstrate the core methodology of the OD variants presented in the main text. Concepts are represented by numbers, and their definitions are given as lists of other concept numbers:
1: [2, 6] 13: [14, 15]
2: [6, 1] 14: [15, 13]
3: [4, 5] 15: [13, 14]
4: [5, 3] 16: [17, 18]
5: [3, 4] 17: [18, 16]
6: [1, 2] 18: [16, 17]
7: [6, 8] 19: [18, 20]
8: [9, 10] 20: [21, 22]
9: [10, 11] 21: [22, 19]
10: [11, 12] 22: [21, 20]
11: [12, 8] 24: [22, 23]
12: [10, 9]
A.2 Example: Weak Ontological Differentiation
This example illustrates the Weak Ontological Differentiation (WOD) process for the initial pair (16, 20) of the sample set, demonstrating the application of the cancelation rule $\kappa_{W}$ and termination rule $\tau_{W}$ given in the main text (see Section II.2 for more information).
As we proceeded with the Strong Ontological Differentiation example described in Table II.2.1 (main text), in the step-by-step illustration below:
  • “Level $n$” refers to the result of the $n$-th recursive application of the read function $R$.
  • The state after each level’s expansion shows the elements generated at all levels up to that point.
  • Canceled elements according to rule $\kappa_{W}$ are marked as ‘(c)’. Cancelation status is updated after each level’s expansion by checking for any element instance that has appeared previously anywhere in the expansion (regardless of the origin side or level).
  • The process terminates when all elements generated at the current level $\omega$ from at least one of the initial concepts (16 or 20) are marked as canceled.
  • The final score is calculated by summing $n\times(\text{count of elements canceled at level }n)$.
Example A.1.
$\text{WOD}(16,20)$:
Level 1: $R^{1}(16)=[17,18]$;$R^{1}(20)=[21,22]$ State after Level 1 expansion: L0: 16 ;  20 L1: 17, 18 ;  21, 22
Level 2: $R^{2}(16)=[18,16,16,17]$;$R^{2}(20)=[22,19,21,20]$ State after Level 2 expansion: L0: 16(c) ;  20(c) L1: 17(c), 18(c) ;  21(c), 22(c) L2: 18(c), 16(c), 16(c), 17(c) ;  22(c), 19, 21(c), 20(c)
Termination: The process terminates at $\omega=2$ according to $\tau_{W}$, because at this level, there is at least one side of at least one level whose elements are all canceled. In this case, both sides at level 0, both sides at level 1, and the left side at level 2 have all of their elements canceled.
[…]
A.3 Example: Great Ontological Differentiation and $\omega_{GOD}$ determination
This example illustrates the Great Ontological Differentiation (GOD) process for the pair (16, 20). GOD primarily differs from WOD and SOD in its termination rule $\tau_{G}$, which depends only on the introduction of new unique elements during the expansion, not on cancelations. The goal here is to find the maximum relevant expansion depth, $\omega_{GOD}$.
Example A.2.
$\text{GOD}(16,20)$:
Elements generated: {16, 20}. All new. Unique elements seen so far: {16, 20}. Level 1: Elements generated: {17, 18, 21, 22}. All new compared to previous levels. Unique elements seen so far: {16, 17, 18, 20, 21, 22}. Level 2: Elements generated: {18, 16, 16, 17, 22, 19, 21, 20}. New unique element: {19}. Unique elements seen so far: {16, 17, 18, 19, 20, 21, 22}. Level 3: Elements generated: {16, 17, 18, 19, 20, 21, 22}. All elements generated ({16, 17, 18, 19, 20, 21, 22}) have been seen in previous levels. No new unique elements. Termination: Stopping at Level 3 because no new elements were found: $\omega_{GOD}=3$.
Appendix B Distributions of SOD and cosine similarity Scores
Figure S1: Distributions for ground-filtered condition. Transformation stages for SOD scores and cosine similarity. (A) Raw SOD scores (note log scale on y-axis). (B) log(1+SOD) scores. (C) Normalized log(1+SOD) scores scaled to $[0,1]$. (D) Raw cosine similarity. (E) log(1+Cosine) scores. (F) Normalized log(1+Cosine) scores scaled to $[0,1]$.
Figure S2: Distributions for random removal condition. Transformation stages for SOD scores and cosine similarity. (A) Raw SOD scores (note log scale on y-axis). (B) log(1+SOD) scores. (C) Normalized log(1+SOD) scores scaled to $[0,1]$. (D) Raw cosine similarity. (E) log(1+Cosine) scores. (F) Normalized log(1+Cosine) scores scaled to $[0,1]$.
Figure S3: Distributions for targeted removal Condition. Transformation stages for SOD scores and cosine similarity. (A) Raw SOD scores (note log scale on y-axis). (B) log(1+SOD) scores. (C) Normalized log(1+SOD) scores scaled to $[0,1]$. (D) Raw cosine similarity. (E) log(1+Cosine) scores. (F) Normalized log(1+Cosine) scores scaled to $[0,1]$.
Appendix C Extended analysis: Network topology and path length
To perform a stricter test of semantic coherence, we extended our analysis to a more structurally constrained directed network, where links strictly follow definitional relationships.
This investigation utilized the same set of 2 million source-target word pairs from the analysis presented in the main text. As in the main results, we analyzed only those pairs where both semantic navigation (SN) and shortest path (SP) algorithms successfully found a path to the target. The number of such valid pairs differs significantly by topology. For the symmetric network, the success rate was consistently high, as we already know, with over 99% of pairs being valid. In contrast, the directed network, being far more constrained, naturally had lower successful pairs: 58,295 for ground-filtered, 19,236 for random removal, and 78,237 for targeted removal.
When applying our original total cumulative SOD metric to this network, we observed a surprising and dramatic reversal of our main finding: the clear advantage held by SN on the fully symmetric network was completely overturned, with SP now appearing superior as shown in Table S1.
Corpus Type SN $<$ SP (%) SN = SP (%) SN $>$ SP (%)
Ground-filtered 15.59 25.05 59.36
Targeted removal 30.20 15.70 54.10
Random removal 13.17 35.69 51.14
Table S1: Comparison of total cumulative SOD on directed networks. In stark contrast to the symmetric network results, SP paths are favored.
The results in Table S1 present a clear contradiction to our findings on the symmetric network. The reversal of SN’s advantage prompted the deeper investigation into the confounding effect of path length detailed in the following sections.
C.1 The confounding effect of path Length
We hypothesized that the discrepancy observed in the directed network could be explained by a systematic difference in the hop count of the paths generated by each algorithm. A detailed breakdown of the average path lengths for both SN and SP across all six experimental conditions is presented in Table S2.
Corpus Type Outcome Mean SN Length Mean SP Length
Directed Network
Ground-filtered OVERALL 8.93 5.11
When SN average SOD $<$ SP 10.05 6.00
When SN average SOD = SP 3.97 3.97
When SN average SOD $>$ SP 11.22 4.91
Random removal OVERALL 8.19 4.63
When SN average SOD $<$ SP 9.44 6.02
When SN average SOD = SP 3.24 3.24
When SN average SOD $>$ SP 12.36 4.82
Targeted removal OVERALL 9.28 6.59
When SN average SOD $<$ SP 9.87 7.84
When SN average SOD = SP 3.55 3.55
When SN average SOD $>$ SP 11.04 6.17
Symmetric Network
Ground-filtered OVERALL 3.40 3.82
When SN average SOD $<$ SP 3.61 4.12
When SN average SOD = SP 2.44 2.44
When SN average SOD $>$ SP 3.57 4.09
Random removal OVERALL 3.63 4.19
When SN average SD $<$ SP 3.80 4.48
When SN average SD = SP 2.69 2.70
When SN average SD $>$ SP 3.79 4.39
Targeted removal OVERALL 3.82 4.27
When SN average SD $<$ SP 3.97 4.55
When SN average SD = SP 2.87 2.87
When SN average SD $>$ SP 4.02 4.49
Table S2: Detailed average path lengths (measured as number of hops). This table shows the average path lengths for SN and SP, both overall and conditioned on the outcome of the average SOD comparison. A clear pattern emerges in the directed network, where SN paths are substantially longer than SP paths.
Before interpreting these results, it is crucial to recall the objective of the SP algorithm as implemented in our study. SP navigation uses Dijkstra’s algorithm to find the path with the minimum possible cumulative cost, which does not always correspond to the path with the fewest hops.
With this clarification in mind, Table S2 reveals a critical insight. In the highly-constrained directed network, the lowest-cost path found by SP is also, as an empirical result, significantly shorter in hop count (e.g., 8.93 vs. 5.11 hops overall for ground-filtered) than the path found by SN. Conversely, in the denser symmetric network, where SP has more freedom to find alternative low-cost routes, the path lengths are much more comparable, with SP paths being just slightly longer than SN paths.
Because the total cumulative SOD score naturally increases with path length, a direct comparison using this metric is clearly confounded in the directed network case. The SP algorithm gains a significant apparent advantage simply by finding paths with fewer, albeit potentially less coherent, steps. This necessitates the use of a standardized metric to make a fair comparison of intrinsic path coherence.
C.2 Standardized analysis: Average SOD per hop
To control for the confounding effect of path length, we introduced a standardized metric: the average SOD per hop, calculated as the total cumulative SOD divided by the number of hops in the path. This metric assesses the semantic ‘cost’ per step, providing a fairer comparison of path coherence regardless of length. Table S3 shows the comparison results using this standardized metric across all six experimental conditions.
Corpus Type SN $<$ SP (%) SN = SP (%) SN $>$ SP (%)
Directed Network
Ground-filtered 40.37 25.05 34.58
Random removal 31.18 35.70 33.11
Targeted removal 49.90 15.70 34.40
Symmetric Network
Ground-filtered 52.46 16.95 30.58
Random removal 54.14 14.80 31.06
Targeted removal 51.57 15.44 32.99
Table S3: Average SOD per hop. Using this standardized metric, which accounts for path length, SN’s advantage is re-established in the symmetric networks (leading to values comparable to those in Table 4, main text) and its performance relative to SP is substantially improved in the directed networks.
Using the average SOD per hop, the results shift significantly, decisively re-establishing the superiority of semantic navigation. The standardization reveals that the individual steps in SN paths are, on average, far more definitionally coherent. In the symmetric network, SN’s advantage becomes even more pronounced, with SN achieving a lower average SOD in a clear majority of cases (winning in over 51–54% of pairs). This trend largely extends to the directed network, where SN paths are now clearly preferable in the ground-filtered case and demonstrate a notable lead in the targeted removal case, with nearly 50% of pairs favoring SN.
However, the directed – random removal condition presents a notable anomaly: the outcome is nearly a three-way tie, with SN having a slightly lower chance of achieving a better score than SP ($\sim$31vs. $\sim$33%). This specific result suggests that while standardization is a major step forward, the profound differences in path length might still be subtly confounding the comparison. To definitively settle the question and isolate the intrinsic coherence of the navigation strategies, we must therefore control for path length completely.
C.3 Controlled comparison: Equal-length paths
The most rigorous method to eliminate the effect of path length is to compare only those pairs where both SN and SP produced paths of the exact same length. This creates a perfectly controlled experimental condition where any difference in the average SOD per hop can be attributed solely to the navigation strategy. Table S4 presents the results for this subset of the data. We also perform a one-sided Wilcoxon signed-rank test to determine if the average SOD for SN paths is statistically significantly lower than that for SP paths.
Corpus Type SN $<$ SP (%) SN = SP (%) SN $>$ SP (%) $p$-value ($H_{\text{a}}$: SN $<$ SP)
Directed Network
Ground-filtered 14.69 81.91 3.40 $6.53\times 10^{-243}$
Random removal 8.68 88.67 2.65 $2.90\times 10^{-40}$
Targeted removal 24.47 68.59 6.94 $<10^{-300}$
Symmetric Network
Ground-filtered 42.34 35.02 22.64 $<10^{-300}$
Random removal 43.77 31.55 24.68 $<10^{-300}$
Targeted removal 42.26 30.34 27.39 $<10^{-300}$
Table S4: Comparison of average SOD for same-length paths. This table shows the outcome percentages and the p-value from a Wilcoxon test on the subset of pairs where both algorithms found paths of equal hop count. The extremely small p-values indicate overwhelming statistical significance, confirming SN’s superior coherence when path length is controlled.
The results in Table S4 are unequivocal. By comparing only paths of identical length, any confounding effect from this variable is eliminated, allowing for a direct assessment of the navigation strategies themselves. A striking feature of the data, particularly for the directed network, is the large proportion of cases where the scores are equal. This is primarily because for many source-target pairs, especially in a constrained, directed topology, both algorithms find the exact same, often unique, optimal path.
The true test of the strategies, therefore, lies in the subset of cases where they produce different paths of the same length. In this decisive subset, the results show a commanding advantage for semantic navigation. For symmetric networks, the ratio of outcomes where SN is preferable (SN $<$ SP) to where SP is preferable (SN $>$ SP) is nearly 2-to-1 (e.g., 42.34% vs. 22.64% for the ground-filtered condition). This advantage becomes even more pronounced for directed networks, where the same ratio approaches 4-to-1 in favor of SN (e.g., 14.69% vs. 3.40% for the ground-filtered condition).
The statistical significance of this finding is confirmed by the one-sided Wilcoxon tests (Table S4), which yield $p$-values that are effectively zero across all six conditions. This provides overwhelming evidence to reject the null hypothesis (i.e., $H_{\text{0}}$: SN $\geq$ SP) and allows us to conclude that, when the effect of path length is removed, SN’s navigation strategy is inherently more faithful to the definitional structure of the lexicon than the strategy employed by SP.
C.4 Statistical modeling of path coherence
C.4.1 Directed network
To statistically model and disentangle the effects of navigation strategy and path length, we first fitted an analysis of covariance (ANCOVA) model exclusively on the data from the three directed network conditions. This provides the most rigorous statistical test of our hypothesis in the most constrained network environment.
Variable Coefficient Std. Error $t$-statistic P $>|t|$
Intercept (SN baseline) $-8.436\times 10^{3}$ $1.304\times 10^{3}$6.467 $<$ 0.001
Path type [SP] $-1.094\times 10^{5}$ $2.069\times 10^{3}$52.850 $<$ 0.001
Length (for SN) $3.223\times 10^{3}$ $1.277\times 10^{2}$ 25.245 $<$ 0.001
Path type [SP] $\times$ Length $\mathbf{2.351\times 10^{4}}$ $\mathbf{2.863\times 10^{2}}$ 82.136 $<$ 0.001
Note: $N=$ 311,536; $R^{2}=$ 0.037
Table S5: ANCOVA results for directed networks. This regression model was fitted on all 311,536 path observations from the three directed network conditions. The model includes an interaction term to test if the effect of path length differs between SN and SP. The reference category for path type is SN.
Figure S4: Semantic coherence vs. path length (directed networks only). Regression plot showing the relationship between path length and the log-transformed average SOD per hop for SN (blue) and SP (red) paths. The plot visualizes the ANCOVA model fitted on all valid paths from the three directed network conditions. The diverging slopes provide a powerful visual confirmation of the significant interaction effect found in our statistical model (Table S5).
The results of directed-only ANCOVA, shown in Table S5 and visualized in Fig. S4, are compelling. The model reveals a crucial dynamic through its interaction term. The slope for SN paths (the length coefficient in Table S5) is relatively shallow ($3.22\times 10^{3}$), indicating a modest increase in definitional divergence per hop. In contrast, the slope for SP paths is the sum of the baseline slope and the large, positive interaction term ($2.35\times 10^{4}$), resulting in a much steeper slope of approximately $2.67\times 10^{4}$.
This demonstrates that while SP may be effective over a single hop, its definitional coherence degrades approximately 8 times faster than SN’s with each subsequent step. This analysis confirms that SN’s strategy results in paths that are fundamentally more coherent over longer distances, even in the most challenging network topology. This robustly supports our central hypothesis.
C.4.2 Global meta-analysis
Finally, to test the universality of this trend, we performed a global meta-analysis by fitting a second ANCOVA model to the entire dataset of 12,286,950 path observations from all six experimental conditions. Because the symmetric network data constitutes over 97% of these observations, this global model primarily serves to confirm that the fundamental relationship observed in the directed networks holds true across the entire experimental landscape, especially in the denser, more highly connected symmetric topology.
Variable Coefficient Std. Error $t$-statistic P $>|t|$
Intercept (SN baseline) $9.288\times 10^{13}$ $1.49\times 10^{12}$ 62.250 $<$ 0.001
Path type [SP] $4.446\times 10^{13}$ $2.50\times 10^{12}$ 17.764 $<$ 0.001
Length (for SN) $-8.500\times 10^{12}$ $3.71\times 10^{11}$22.900 $<$ 0.001
Path type [SP] $\times$ Length $\mathbf{1.260\times 10^{13}}$ $\mathbf{5.97\times 10^{11}}$ 21.084 $<$ 0.001
Note: $N=$ 12,286,950; $R^{2}=$0.001
Table S6: ANCOVA meta-analysis of average SOD per hop. This regression model was fitted on all 12,286,950 path observations from the six experimental conditions. The model includes an interaction term to test if the effect of path length differs between SN and SP. The reference category for path type is SN.
Figure S5: Visualizing the ANCOVA meta-analysis: Semantic coherence vs. path length. Regression plot showing the relationship between path length and the log-transformed average SOD per hop for semantic navigation (SN, blue) and shortest path (SP, red) paths. The plot is generated from a random sample of 100,000 data points from all six experimental conditions for visual clarity, but the trend lines represent the linear regression fits from the ANCOVA model on the full dataset of over 12 million observations (Table S6). The diverging slopes provide a powerful visual confirmation of the significant interaction effect found in our statistical model. The nearly flat slope for SN demonstrates that its per-step semantic coherence is remarkably stable regardless of path length. In contrast, the positive slope for SP shows that its coherence degrades as paths become longer. This illustrates that SN maintains its semantic coherence over longer paths more effectively than SP, and its advantage grows with distance. The wide scatter of points also visually represents the high intrinsic variance in the data, justifying the modest $R^{2}$ value of the model.
The global model, summarized in Table S6, corroborates the findings from the directed-only analysis on a much larger scale. While the $R^{2}$ value is modest, which is expected in a dataset of this scale with high intrinsic variance, the statistical significance of the coefficients provides robust evidence for the direction and relative importance of these effects. The model allows us to dissect the relationship between path type, path length, and semantic coherence.
The key coefficients of the regression model are all highly significant ($p<0.001$), leading to a clear interpretation:
  1. 1.
    A fundamental difference: the coefficient for Path type [SP] ($4.446\times 10^{13}$) represents the baseline difference between the two strategies. It shows that even at a theoretical path length of zero, an SP path is predicted to have a significantly higher average SOD. This establishes that SN is fundamentally more coherent from the very first step.
  2. 2.
    Diverging trajectories: The most critical result is the significant, positive interaction term, Path type [SP] $\times$ Length ($1.260\times 10^{13}$). This term proves that the two strategies do not just have a constant difference; their performance trajectories diverge as paths become longer.
    • The slope for Length (for SN) is negative ($-8.500\times 10^{12}$), indicating that SN paths maintain or even improve their per-step coherence over longer distances.
    • The slope for SP paths is the sum of the main length effect and the interaction term, resulting in a positive slope ($4.1\times 10^{12}$). This demonstrates that the semantic coherence of SP paths actively degrades with each additional step.
In essence, after accounting for all variations from network type and corpus processing, the model reveals a universal trend: SN paths are not only inherently more coherent than SP paths, but this advantage in coherence grows with path length.
C.5 Conclusion
In summary, this extensive robustness analysis confirms that the central conclusion of our paper holds true with remarkable consistency. Across all tested network topologies and corpus processing schemes, after controlling for path length, semantic navigation consistently generates paths with significantly lower definitional divergence than the shortest-path algorithm.
The controlled comparisons of same-length paths and the global ANCOVA model both provide overwhelming evidence for this conclusion. This validates Ontological Differentiation as a sensitive and powerful metric for path coherence and underscores the universal effectiveness of semantic vector-based navigation in producing coherent trajectories through lexical networks.